# MARFT: Multi-Agent Reinforcement Fine-Tuning

### v2

@@ Abstract
Large Language Model (LLM)-based Multi-Agent Systems (LaMAS) have demonstrated remarkable capabilities in addressing complex, agentic tasks requiring multifaceted reasoning and collaboration, from generating high-quality presentation slides to conducting sophisticated scientific research. Meanwhile, Reinforcement Learning (RL) has been widely recognized for its effectiveness in enhancing agent intelligence, but limited research has investigated the fine-tuning of LaMAS using foundational RL techniques. Moreover, the direct application of conventional Multi-Agent Reinforcement Learning (MARL) methodologies to LaMAS introduces significant challenges, stemming from the unique characteristics and mechanisms inherent to LaMAS. To address these challenges, this article presents a comprehensive study of LLM-based MARL and proposes a novel paradigm termed Multi-Agent Reinforcement Fine-Tuning (MARFT). We introduce a universal algorithmic framework tailored specifically for LaMAS, outlining the conceptual foundations, key distinctions, and practical implementation strategies. We begin by reviewing the evolution from traditional RL to Reinforcement Fine-Tuning (RFT), setting the stage for a parallel analysis in the multi-agent domain. In the context of LaMAS, we elucidate critical differences between classical MARL and MARFT, such as asynchronous agent interactions, profile-aware agent design, and heterogeneous architectural configurations, etc. These differences motivate a transition toward a novel, LaMAS-oriented formulation of RFT. Central to this work is the presentation of a robust and scalable MARFT framework. We detail the core algorithm, emphasizing its modularity and adaptability, and provide a complete, open-source implementation to facilitate adoption and further research. The latter sections of the paper explore real-world application perspectives and opening challenges in MARFT, including dynamic environment modeling, sample inefficiency, and the current lack of cohesive frameworks. By bridging theoretical underpinnings with practical methodologies, this work aims to serve as a roadmap for researchers seeking to advance MARFT toward resilient, adaptive, and human-aligned solutions in agentic systems. Our implementation of the proposed framework is publicly available at: https://github.com/jwliao-ai/MARFT.
@@ 1 Introduction
Multi-Agent Systems (MAS) have consistently demonstrated superior capabilities in addressing complex agentic tasks relative to their single-agent counterparts. This is evident in the latest General AI Assistants (GAIA) leaderboard (Mialon et al., 2024), where the top-performing systems are all multi-agent frameworks. Prominent frameworks, including OpenAI Agents SDK (formerly Swarm), Microsoft AutoGen (Wu et al., 2023), Magnetic-One, CAMEL-AI OWL (CAMEL-AI.org, 2025), and Google’s AI Co-Scientist (Gottweis et al., 2025), further reinforce the growing importance of multi-agent architectures. Notably, in OpenAI’s vision for AGI, multi-agent systems constitute the highest organizational layer. However, as Cemri et al. (2025) illustrates, multi-agent systems are prone to a variety of failure modes—fourteen in total—related to system design, coordination, and control, where MARL is poised to address these challenges. Besides, recently Yang et al. (2025b) proposed the first framework to systematically quantify the contributions of different modules in multi-agent systems, potentially illuminating the direction for future multi-agent optimization in LLMs.
@@ 2.1.2 Reinforcement Learning (RL)
The Objective of RL. In classic reinforcement learning settings, an agent is capable of perceiving its environment, making decisions, and taking actions, while learning from past experiences to enhance its performance. The agent interacts with the environment until certain terminal conditions are met, generating a trajectory $\tau=(s_{0},a_{0},r_{0},s_{1},a_{1},r_{1},\cdots,s_{T},a_{T},r_{T})$ of length $T$, with a cumulative reward expressed as $g=\sum_{t=0}^{T}\gamma^{t}r_{t}$. The primary objective is for the agent to achieve a high cumulative reward in a dynamic and uncertain environment. This is precisely what RL aims to accomplish—obtaining a policy that enables the agent to dynamically act in dynamic environments to maximize cumulative reward. Mathematically, the objective is formulated as $\max\limits_{\pi}\mathbb{E}_{\tau\sim p_{\pi}(\tau)}\left[\sum_{t=0}^{T}\gamma^{t}R_{t}\right]$ (Sutton & Barto, 1998). Here, the cumulative reward from timestep $t$ is denoted as $G_{t}=\sum_{t^{\prime}=t}^{\infty}\gamma^{t^{\prime}-t}R_{t^{\prime}}$.
Value-based Methods. Given that RL optimizes the policy by leveraging rewards to incentivize the policy to learn actions that yield the highest rewards, it is natural to consider a value-based approach. This value-based approach provides an elegant and effective solution to the critical credit assignment problem (Minsky, 1961; Sutton & Barto, 1998). This approach focuses on learning indirect rewards, i.e, values, which estimate the goodness of an action before receiving the reward and ideally can better optimize the policy in long-horizon and reward-sparse tasks. To this end, two vital proxies to the RL objective are introduced: the state-action value function (Q-function) and the state value function (V-function) (Sutton & Barto, 1998). Their definitions are presented respectively as
@@ 8             $\mathcal{D}\leftarrow\mathcal{D}\cup\{\left(s,a,s^{\prime},r\right)\}$
Policy-based Methods. Another approach to optimizing policies in RL involves directly estimating the gradient of the policy and performing gradient ascent. This method is largely based on the policy gradient (PG) theorem (Sutton et al., 1999). Recall that the objective is to maximize the expected reward: $J(\theta)=\mathbb{E}_{{}_{\tau\sim p_{\pi_{\theta}}(\tau)}}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right]$. [^1] In other papers you may see people denote it $J(\pi_{\theta})$ In fact, these two expressions are the same. You can view it as some reward $J$ being a function of parameters $\theta$ or some reward $J$ being a function of the policy $\pi$, which is parameterized by $\theta$. According to the policy gradient theorem, the gradient of the expected reward $J(\theta)$ can be expressed as
@@
Actor-Critic. Actor-Critic (AC) (Sutton & Barto, 1998; Konda & Tsitsiklis, 1999; Degris et al., 2012; Mnih et al., 2016) architecture is a hybrid way, incorporating both value-based methods and policy-based methods, to optimize the policy. While value-based methods extract policy poorly and policy-based methods lack direct estimations to values with a result of low efficiency, Actor-Critic inherits fine-grained credit assignment from value-based methods and optimization stability from policy-based methods. Actor-Critic algorithms train the value function and policy iteratively. For each iteration, it updates the value function based on TD error and optimizes the policy via the policy gradient method
@@ 2.1.3 From RL to RFT
Differences. Though RL and RFT share foundational methodological principles, there still exist nuanced differences as listed in Table 1. RL, closely rooted in optimal control frameworks, typically trains policy agents, such as robots, from scratch, requiring no prior domain knowledge. In contrast, RFT operates on large-scale pretrained models and adapts these models to specialized tasks. Additionally, RL trains a policy specifically for a particular task within a specific environment, while RFT enhances a model’s task-specific capabilities without degrading its original abilities, such as language understanding or generative skills in LLMs. The environments also differ significantly. Conventional RL environments are entirely stochastic, whereas typical RFT agentic environments for LLMs can integrate both deterministic and stochastic transitions—deterministic for operations like sentence or token concatenation, and stochastic for environment feedback, such as tool execution results or next user queries. Since LLMs generate tokens to form sentences, RFT can operate at both the action-level (sentence-level) and token-level, whereas traditional RL settings are typically action-level. Another major difference is that RL is known for its low sample efficiency, while RFT requires relatively fewer, high-quality samples.
@@
Process Reset state, then sample traj. Reset query, sample traj., then check answer
@@
Table 1: Differences between RL and RFT.
Implementations. Since we know the difference between RFT and RL, we can introduce lots of fine-tuning techniques to better show how RFT works and what is reflected in many existing research. The first is LoRA, which freezes the base model and injects trainable rank decomposition matrices into the model. LoRA is vital for fine-tuning LLMs, especially for LaMAS, because it can optimize the policy efficiently with minimal parameter updates and also makes LLM-based multi-agent training affordable and ideal. As a single model can have different LoRA adapters for different tasks, multi-agent systems can be made using only one base model equipped with multiple adapters (2024). Secondly, it becomes necessary to implement divergence constraints during training to prevent the policy update from straying too far from the initial or reference policy, which possesses the pretrained abilities. PPO and other trust-region-related algorithms provide strong guarantees for monotonic improvement, and in their objectives, we can also modify the KL divergence from between $\pi_{\text{old}}$ to between $\pi_{\text{ref}}$ as GRPO (Shao et al., 2024) does in (13).
Prevailing Objective Functions. To further elaborate on the objective functions used in RFT, we will introduce several prominent methods, including GRPO, DAPO, and VAPO. These objective functions are designed to optimize the policy updates while ensuring stability and efficiency during the fine-tuning process.
@@
$\text{where}\ L_{\text{NLL}}(\theta)-\frac{\sum_{o_{i}\in\mathcal{T}}\sum_{t=1}^{|o_{i}|}\log\pi_{\theta}(a_{t}|s_{t})}{\sum_{o_{i}\in\mathcal{T}}|o_{i}|}.$
@@ 2.2 LLM-based Multi-Agent Systems (LaMAS)
  • Parameter fine-tuning. A more sophisticated approach involves updating agent parameters or network configurations to optimize performance. This paradigm integrates deep learning and specific methodologies, such as multi-agent debating to generate high-quality training datasets (Subramaniam et al., 2025) or fine-tuning programming modules for improved workflow organization (Khattab et al., 2024). Ma et al. (2024) duplicates the LLM into two agents, i.e., a pioneer and an observer, that interact through knowledge transfer and role exchange, improving training stability and performance. However, many parameter fine-tuning methods lack a theoretical foundation in RL, relying instead on ad-hoc techniques. This article will focus on RFT methodologies in subsequent sections, providing a principled framework for LaMAS optimization.
@@ 2.3.1 Decentralized Partially Observable Markov Decision Process
MARL is a variant of RL, which including multiple agents acting simultaneously once receiving the environment observations, and it’s often modeled by a certain form of MDP – Decentralized POMDP (DEC-POMDP) (Oliehoek & Amato, 2016) $\langle\mathcal{N},{\mathcal{S}},\bm{\mathcal{O}},\bm{\mathcal{A}},\mathcal{T},\mathcal{R},\gamma\rangle$, where $\mathcal{N}=\{1,\dots,n\}$ is the set of agents without any order constraint, ${\mathcal{S}}$ is the state space, $\bm{\mathcal{O}}=\prod^{n}_{i=1}\mathcal{O}^{i}$ is the joint observation space, which is a product of local observation spaces at global state $s$, $\bm{\mathcal{A}}=\prod^{n}_{i=1}\mathcal{A}^{i}$ is the joint action space, which is similarly a product of local action spaces, and for the others, we can borrow the explanation from Section 2.1.1.
@@ 2.3.2 Main Approaches of MARL
Independent Learning. One of the most intuitive approaches to implementing MARL is through independent learning. In this setting, each agent operates as an independent entity, possessing its own policy and value function. Other agents are treated as part of the environment, and there is no direct communication between agents. Consequently, each agent optimizes its actions based solely on its own observations and rewards, without considering the policies of other agents.
@@
Centralized Training with Decentralized Execution. Apart from independent learning, another way to MARL is by centralized training. During training, agents share a joint global value function $Q^{\bm{\pi}}(\mathbf{o},a^{1},\dots,a^{n})$ which takes into account the environment state and actions of other agents. For its design, $Q$ can either output the value for the joint action and marginalize out for a specific action, as in COMA (Foerster et al., 2018), or directly output the value for action $a^{i}$ by agent $i$, as in MADDPG (Lowe et al., 2017). The latter approach allows for arbitrary reward structures for separate agents. Once the value learning component is established, policy optimization can be performed using the policy gradient: $\nabla_{\theta^{i}}J(\theta^{i})=\mathbb{E}_{s\sim p^{\bm{\mu}},a^{i}\sim\bm{\pi}^{i}}\left[\nabla_{\theta^{i}}\log\bm{\pi}^{i}(a^{i}|o^{i})Q^{\bm{\pi}}(\mathbf{o},a^{1},\dots,a^{n})\right]$ to update the policies.
As mentioned in Section 2.1.2, PPO has consistently been one of the most effective RL methods. Fortunately, it also extends well to multi-agent scenarios, leading to Multi-Agent PPO (MAPPO) (Yu et al., 2022). Then, the objective of MAPPO can is derived as
@@
Communication and Coordination. Another intriguing approach to MARL involves directly learning the communication among agents. For instance, designing a communication module or gate module can help establish an effective explicit communication and coordination mechanism within the multi-agent system.
@@
Sequential Modeling. Another intriguing perspective on understanding MARL is through sequential modeling (SM). The connection between MARL and SM is grounded in the Multi-Agent Action Decomposition Theorem (Kuba et al., 2021), which is expressed as
$A^{\bm{\pi}}\left(\bm{o},\bm{a}^{1:n}\right)=\sum_{i=1}^{n}A^{\bm{\pi}}\left(\bm{o},\bm{a}^{1:i-1},a^{i}\right).$ (17)
@@
$L(\phi)=\frac{1}{nT}\sum_{i=1}^{n}\sum_{t=0}^{T-1}\left[r_{t}(\bm{o}_{t},\bm{a}_{t})+\gamma V_{\phi}(\hat{\bm{o}}_{t+1}^{i})-V_{\phi}(\hat{\bm{o}}_{t}^{i})\right]^{2}.$ (18)
But for the policy optimization, the modified clipping PPO objective goes as
@@
Networked Agents. One more novel cooperative MARL paradigm involves networked agents, which operate in fully decentralized settings. In this framework, agents perceive local environmental observations and synchronize information locally with neighboring agents over a network (Zhang et al., 2018). A consensus mechanism has been introduced to the paradigm, expanding its applicability to real-world domains with strict privacy requirements (Chen et al., 2022; Varela et al., 2025). By enabling localized information sharing and coordination, networked agents enhance system robustness and adaptability while preserving data privacy, making this paradigm particularly promising for applications in decentralized and privacy-sensitive environments, especially for LaMAS in use.

### v3

@@ Abstract
Large Language Model (LLM)-based Multi-Agent Systems (LaMAS) have demonstrated remarkable capabilities in addressing complex, agentic tasks requiring multifaceted reasoning and collaboration, from generating high-quality presentation slides to conducting sophisticated scientific research. Meanwhile, Reinforcement Learning (RL) has been widely recognized for its effectiveness in enhancing agent intelligence, but limited research has investigated the fine-tuning of LaMAS using foundational RL techniques. Moreover, the direct application of conventional Multi-Agent Reinforcement Learning (MARL) methodologies to LaMAS introduces significant challenges, stemming from the unique characteristics and mechanisms inherent to LaMAS. To address these challenges, this article presents a comprehensive study of LLM-based MARL and proposes a novel paradigm termed Multi-Agent Reinforcement Fine-Tuning (MARFT). We introduce a brand-new POMDP called Flex-POMDP, which aligns with the LaMAS optimization in real world applications and a universal algorithmic framework tailored specifically for LaMAS, outlining the conceptual foundations, key distinctions, and practical implementation strategies. We begin by reviewing the evolution from traditional RL to Reinforcement Fine-Tuning (RFT), setting the stage for a parallel analysis in the multi-agent domain. In the context of LaMAS, we elucidate critical differences between classical MARL and MARFT, such as asynchronous agent interactions, profile-aware agent design, and heterogeneous architectural configurations, etc. These differences motivate a transition toward a novel, LaMAS-oriented formulation of RFT. Central to this work is the presentation of a robust and scalable MARFT framework. We detail the core algorithm, emphasizing its modularity and adaptability, and provide a complete, open-source implementation to facilitate adoption and further research. The latter sections of the paper explore real-world application perspectives and opening challenges in MARFT, including dynamic environment modeling, sample inefficiency, and the current lack of cohesive frameworks. By bridging theoretical underpinnings with practical methodologies, this work aims to serve as a roadmap for researchers seeking to advance MARFT toward resilient, adaptive, and human-aligned solutions in agentic systems. Our implementation of the proposed framework is publicly available at: https://github.com/jwliao-ai/MARFT.
@@ 1 Introduction
Multi-Agent Systems (MAS) have consistently demonstrated superior capabilities in addressing complex agentic tasks relative to their single-agent counterparts. This is evident in the latest General AI Assistants (GAIA) leaderboard (Mialon et al., 2024), where the top-performing systems are all multi-agent frameworks. Prominent frameworks, including OpenAI Agents SDK (formerly Swarm), Microsoft AutoGen (Wu et al., 2023), Magnetic-One, CAMEL-AI OWL (CAMEL-AI.org, 2025), and Google’s AI Co-Scientist (Gottweis et al., 2025), further reinforce the growing importance of multi-agent architectures. Notably, in OpenAI’s vision for AGI, multi-agent systems constitute the highest organizational layer. However, as Cemri et al. (2025) illustrates, multi-agent systems are prone to a variety of failure modes—fourteen in total—related to system design, coordination, and control, where MARL is poised to address these challenges. Besides, recently Yang et al. (2025c) proposed the first framework to systematically quantify the contributions of different modules in multi-agent systems, potentially illuminating the direction for future multi-agent optimization in LLMs.
@@ 2.1.2 Reinforcement Learning (RL)
The Objective of RL
In classic reinforcement learning settings, an agent is capable of perceiving its environment, making decisions, and taking actions, while learning from past experiences to enhance its performance. The agent interacts with the environment until certain terminal conditions are met, generating a trajectory $\tau=(s_{0},a_{0},r_{0},s_{1},a_{1},r_{1},\cdots,s_{T},a_{T},r_{T})$ of length $T$, with a cumulative reward expressed as $g=\sum_{t=0}^{T}\gamma^{t}r_{t}$. The primary objective is for the agent to achieve a high cumulative reward in a dynamic and uncertain environment. This is precisely what RL aims to accomplish—obtaining a policy that enables the agent to dynamically act in dynamic environments to maximize cumulative reward. Mathematically, the objective is formulated as $\max\limits_{\pi}\mathbb{E}_{\tau\sim p_{\pi}(\tau)}\left[\sum_{t=0}^{T}\gamma^{t}R_{t}\right]$ (Sutton & Barto, 1998). Here, the cumulative reward from timestep $t$ is denoted as $G_{t}=\sum_{t^{\prime}=t}^{\infty}\gamma^{t^{\prime}-t}R_{t^{\prime}}$.
Value-based Methods
Given that RL optimizes the policy by leveraging rewards to incentivize the policy to learn actions that yield the highest rewards, it is natural to consider a value-based approach. This value-based approach provides an elegant and effective solution to the critical credit assignment problem (Minsky, 1961; Sutton & Barto, 1998). This approach focuses on learning indirect rewards, i.e, values, which estimate the goodness of an action before receiving the reward and ideally can better optimize the policy in long-horizon and reward-sparse tasks. To this end, two vital proxies to the RL objective are introduced: the state-action value function (Q-function) and the state value function (V-function) (Sutton & Barto, 1998). Their definitions are presented respectively as
@@ 8             $\mathcal{D}\leftarrow\mathcal{D}\cup\{\left(s,a,s^{\prime},r\right)\}$
Policy-based Methods
Another approach to optimizing policies in RL involves directly estimating the gradient of the policy and performing gradient ascent. This method is largely based on the policy gradient (PG) theorem (Sutton et al., 1999). Recall that the objective is to maximize the expected reward: $J(\theta)=\mathbb{E}_{{}_{\tau\sim p_{\pi_{\theta}}(\tau)}}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right]$. [^1] In other papers you may see people denote it $J(\pi_{\theta})$ In fact, these two expressions are the same. You can view it as some reward $J$ being a function of parameters $\theta$ or some reward $J$ being a function of the policy $\pi$, which is parameterized by $\theta$. According to the policy gradient theorem, the gradient of the expected reward $J(\theta)$ can be expressed as
@@
Actor-Critic
Actor-Critic (AC) (Sutton & Barto, 1998; Konda & Tsitsiklis, 1999; Degris et al., 2012; Mnih et al., 2016) architecture is a hybrid way, incorporating both value-based methods and policy-based methods, to optimize the policy. While value-based methods extract policy poorly and policy-based methods lack direct estimations to values with a result of low efficiency, Actor-Critic inherits fine-grained credit assignment from value-based methods and optimization stability from policy-based methods. Actor-Critic algorithms train the value function and policy iteratively. For each iteration, it updates the value function based on TD error and optimizes the policy via the policy gradient method
@@ 2.1.3 From RL to RFT
Table 1: Differences between RL and RFT.
@@
Process Reset state and sample traj. Reset query, sample traj. and verify answer
@@
Differences
Though RL and RFT share foundational methodological principles, there still exist nuanced differences as listed in Table 1. RL, closely rooted in optimal control frameworks, typically trains policy agents, such as robots, from scratch, requiring no prior domain knowledge. In contrast, RFT operates on large-scale pretrained models and adapts these models to specialized tasks. Additionally, RL trains a policy specifically for a particular task within a specific environment, while RFT enhances a model’s task-specific capabilities without degrading its original abilities, such as language understanding or generative skills in LLMs. The environments also differ significantly. Conventional RL environments are entirely stochastic, whereas typical RFT agentic environments for LLMs can integrate both deterministic and stochastic transitions—deterministic for operations like sentence or token concatenation, and stochastic for environment feedback, such as tool execution results or next user queries. Since LLMs generate tokens to form sentences, RFT can operate at both the action-level (sentence-level) and token-level, whereas traditional RL settings are typically action-level. Another major difference is that RL is known for its low sample efficiency, while RFT requires relatively fewer, high-quality samples.
Implementations
Since we know the difference between RFT and RL, we can introduce lots of fine-tuning techniques to better show how RFT works and what is reflected in many existing research. The first is LoRA, which freezes the base model and injects trainable rank decomposition matrices into the model. LoRA is vital for fine-tuning LLMs, especially for LaMAS, because it can optimize the policy efficiently with minimal parameter updates and also makes LLM-based multi-agent training affordable and ideal. As a single model can have different LoRA adapters for different tasks, multi-agent systems can be made using only one base model equipped with multiple adapters (2024). Secondly, it becomes necessary to implement divergence constraints during training to prevent the policy update from straying too far from the initial or reference policy, which possesses the pretrained abilities. PPO and other trust-region-related algorithms provide strong guarantees for monotonic improvement, and in their objectives, we can also modify the KL divergence from between $\pi_{\text{old}}$ to between $\pi_{\text{ref}}$ as GRPO (Shao et al., 2024) does in (13).
Prevailing Objective Functions
To further elaborate on the objective functions used in RFT, we will introduce several prominent methods, including GRPO, DAPO, and VAPO. These objective functions are designed to optimize the policy updates while ensuring stability and efficiency during the fine-tuning process.
@@
$\text{where}\ L_{\text{NLL}}(\theta)=-\frac{\sum_{o_{i}\in\mathcal{T}}\sum_{t=1}^{|o_{i}|}\log\pi_{\theta}(a_{t}|s_{t})}{\sum_{o_{i}\in\mathcal{T}}|o_{i}|}.$
@@ 2.2 LLM-based Multi-Agent Systems (LaMAS)
  • Parameter fine-tuning. A more sophisticated approach involves updating agent parameters or network configurations to optimize performance. This paradigm integrates deep learning and specific methodologies, such as multi-agent debating to generate high-quality training datasets (Subramaniam et al., 2025) or fine-tuning programming modules for improved workflow organization (Khattab et al., 2024). CORY (Ma et al., 2024) duplicates the LLM into two agents, i.e., a pioneer and an observer, that interact through knowledge transfer and role exchange, improving training stability and performance. However, many parameter fine-tuning methods lack a theoretical foundation in RL, relying instead on ad-hoc techniques. SiriuS (Zhao et al., 2025), a self-improving, multi-agent optimization framework using bootstrapped reasoning, iteratively refines agent behaviors by leveraging successful reasoning trajectories and augmenting unsuccessful ones. The most related work, MAPoRL (Park et al., 2025), rooted in MARL, pioneers in applying multi-agent PPO to post-train LLMs, but does not consider the unique characteristics of LaMAS and assumes homogeneous agent behavior through collaborative debate, diverging from real-world scenarios where agents often perform diverse or orthogonal roles to solve complex problems. This article will focus on RFT and MARL-related methodologies in subsequent sections, providing a principled paradigm for general LaMAS reinforcement.
@@ 2.3.1 Decentralized Partially Observable Markov Decision Process
MARL is a variant of RL, which includes multiple agents acting simultaneously once receiving the environment observations, and it’s often modeled by a certain form of MDP – Decentralized POMDP (DEC-POMDP) (Oliehoek & Amato, 2016) $\langle\mathcal{N},{\mathcal{S}},\bm{\mathcal{O}},\bm{\mathcal{A}},\mathcal{T},\mathcal{R},\gamma\rangle$, where $\mathcal{N}=\{1,\dots,n\}$ is the set of agents without any order constraint, ${\mathcal{S}}$ is the state space, $\bm{\mathcal{O}}=\prod^{n}_{i=1}\mathcal{O}^{i}$ is the joint observation space, which is a product of local observation spaces at global state $s$, $\bm{\mathcal{A}}=\prod^{n}_{i=1}\mathcal{A}^{i}$ is the joint action space, which is similarly a product of local action spaces, and for the others, we can borrow the explanation from Section 2.1.1.
@@ 2.3.2 Main Approaches of MARL
Independent Learning
One of the most intuitive approaches to implementing MARL is through independent learning. In this setting, each agent operates as an independent entity, possessing its own policy and value function. Other agents are treated as part of the environment, and there is no direct communication between agents. Consequently, each agent optimizes its actions based solely on its own observations and rewards, without considering the policies of other agents.
@@
Centralized Training with Decentralized Execution
Apart from independent learning, another way to MARL is by centralized training. During training, agents share a joint global value function $Q^{\bm{\pi}}(\mathbf{o},a^{1},\dots,a^{n})$ which takes into account the environment state and actions of other agents. For its design, $Q$ can either output the value for the joint action and marginalize out for a specific action, as in COMA (Foerster et al., 2018), or directly output the value for action $a^{i}$ by agent $i$, as in MADDPG (Lowe et al., 2017). The latter approach allows for arbitrary reward structures for separate agents. Once the value learning component is established, policy optimization can be performed using the policy gradient: $\nabla_{\theta^{i}}J(\theta^{i})=\mathbb{E}_{s\sim p^{\bm{\mu}},a^{i}\sim\bm{\pi}^{i}}\left[\nabla_{\theta^{i}}\log\bm{\pi}^{i}(a^{i}|o^{i})Q^{\bm{\pi}}(\mathbf{o},a^{1},\dots,a^{n})\right]$ to update the policies.
As mentioned in Section 2.1.2, PPO has consistently been one of the most effective RL methods. Fortunately, it also extends well to multi-agent scenarios, leading to Multi-Agent PPO (MAPPO) (Yu et al., 2022). Then, the objective of MAPPO is derived as
@@
Communication and Coordination
Another intriguing approach to MARL involves directly learning the communication among agents. For instance, designing a communication module or gate module can help establish an effective explicit communication and coordination mechanism within the multi-agent system.
@@
Sequential Modeling
Another intriguing perspective on understanding MARL is through sequential modeling (SM). The connection between MARL and SM is grounded in the Multi-Agent Advantage Decomposition Theorem (Kuba et al., 2021) as below:
Theorem 1
(Multi-Agent Advantage Decomposition Theorem (Kuba et al., 2021)). For any predefined permutation of $n$ agents, for any joint observation $\bm{o}\in\bm{\mathcal{O}}$ and joint action $\bm{a}\in\bm{\mathcal{A}}$, the following always holds:
$A^{\bm{\pi}}(\bm{o},\bm{a}^{1:n})=\sum_{m=1}^{n}A^{\bm{\pi}}(\bm{o},\bm{a}^{1:m-1},\bm{a}^{m}).$ (17)
@@
$L(\phi)=\frac{1}{nT}\sum_{i=1}^{n}\sum_{t=0}^{T-1}\left[r_{t}(\bm{o}_{t},\bm{a}_{t})+\gamma V_{\bar{\phi}}(\hat{\bm{o}}_{t+1}^{i})-V_{\phi}(\hat{\bm{o}}_{t}^{i})\right]^{2},$ (18)
where $\bar{\phi}$ is the parameters of the target network, which are non-differentiable and updated at intervals. But for the policy optimization, the modified clipping PPO objective goes as
@@
Networked Agents
One more novel cooperative MARL paradigm involves networked agents, which operate in fully decentralized settings. In this framework, agents perceive local environmental observations and synchronize information locally with neighboring agents over a network (Zhang et al., 2018). A consensus mechanism has been introduced to the paradigm, expanding its applicability to real-world domains with strict privacy requirements (Chen et al., 2022; Varela et al., 2025). By enabling localized information sharing and coordination, networked agents enhance system robustness and adaptability while preserving data privacy, making this paradigm particularly promising for applications in decentralized and privacy-sensitive environments, especially for LaMAS in use.